Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability!

Expert replies
by Ozlemg » Tue Jul 05, 2011 4:54 am
A box contains ten assemblies of which two are defective. A sample of three assemblies is selected at random. What is the probability that the two defective parts will be selected?

A.0.0667
B.0.0776
C.0.667
D.0.433
E.0.20

Why is my colculation not correct? Where am I wrong?
Probability of Defective : 0.2
Probability of Non Defective : 0.8

Selecting 2 defective ones:
1/5*1/5*4/5 =4/125

OA is A
The more you suffer before the test, the less you will do so in the test! :)
Join the discussion
Source: — Problem Solving |

by Frankenstein » Tue Jul 05, 2011 5:10 am
Hi,
Both the defensive ones are to be picked.This can be done in 1 way
From the remaining 8 one has to be picked in 8C1 ways.
Total number of ways of picked 3 out of 10 is 10C3
So, probability = 1*8C1/10C3 = 8/120 = 1/15

Hence, A

Coming to your method.
1st defective is selected in 2/10 ways
After it is picked 2nd one is picked in 1/9 ways
Third one is picked in 8/8 ways
So, probability is 1/5*1/9*1 = 1/45
But, here we are considering the order. In this case, the non-defective one is picked 3rd. It can be picked first or second as well
When non-defective one is picked first, probability is (8/10)*(2/9)*(1/8) = 1/45
When non-defective one is picked second, probability is (2/10)*(8/9)*(1/8) = 1/45
So, probability is 3(1/45) = 1/15.
Cheers!

Things are not what they appear to be... nor are they otherwise
Join the discussion

by winniethepooh » Tue Jul 05, 2011 5:14 am
Answer is A.
The answer should be solved in this way:
You have in all 10C3 choices that is 120 choices, so this is your denominator.
In the Numerator you have 2 fixed choices to get the 2 defectives so 2C2 * 8C1(which is any assembly out of the remaining 8)
Hence, 8/120 = 0.0667
Join the discussion